If a tennis ball container has a diameter of 2.625 inches and a height of 7.875 inches how much wasted space is there in the can (3 tennis balls) if 1 tennis ball has a diameter of 2.5 inches? I think you use the formula for volume of a cylinder and volume of sphere, but I'm getting lost. Thanks so much, Cathy
Answer by Alan3354(69443)

If a tennis ball container has a diameter of 2.625 inches and a height of 7.875 inches how much wasted space is there in the can (3 tennis balls) if 1 tennis ball has a diameter of 2.5 inches?
-------------------
Vol of the can = pi*r^2*h
VC = pi*(2.625/2)^2*7.875
VC =~ 13.5659*pi cubic inches
--------------
Vol of 1 ball = (4/3)*pi*r^3
VB = (4/3)*pi*1.25^3
VB =~ 2.604*pi cubic inches
VB3 = 7.8125*pi cubic inches
----------------
VC - VB3 =~5.7534*pi in^3
=~ 18.075 in^2